OpenStudy (jellybot23):

Compare the equation for a line through the origin, y = mx, to the equation Fc= mv^2/r , and explain what the slope of the graph of Fc vs. v2 represents. Remember that the m in the equation for a line represents the slope, and the m in the centripetal force equation represents the mass of the stopper. I have no idea what they are asking with comparing the slopes

OpenStudy (jellybot23):

@Luigi0210

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle y=mx }\) A line with a positive slope \(\color{#000000 }{ \displaystyle m }\), that goes through the origin. Sometimes, this would be referred to as, \(\tt"\)direct variation\(\tt"\). The \(\color{#000000 }{ \displaystyle x }\) is proportional to \(\color{#000000 }{ \displaystyle y }\), and the proportion constant is \(\color{#000000 }{ \displaystyle m }\). \(\color{#000000 }{ \rm F_c=\frac{\large m\times v^2}{\large r} }\) The equation of the Cetripetal Force, of some particle that moves along a circle. where, \(\color{#000000 }{ \displaystyle {\rm F_c} }\) Centripetal Force \(\color{#000000 }{ \displaystyle \rm m }\) Mass \(\color{#000000 }{ \displaystyle \rm v }\) Velocity of the partical \(\color{#000000 }{ \displaystyle \rm r }\) Radius of the circle \(\color{#000000 }{ \displaystyle \rm A_{\large c} }\) Centripetal acceleration The cerntripetal acceleration is given by \(\color{#000000 }{ \rm A_{\large c}=\frac{\Large v^2}{\Large r} }\) So, alternatively, you can write \(\color{#000000 }{ \rm F_{\large c}=m\times A_{\large c} }\) And that is also a line that shows proportion between cetripetal force and centripetal acceleration. (This force equation is just like any other \(F=ma\))